# VTG-LLM: Integrating Timestamp Knowledge into Video LLMs for Enhanced Video Temporal Grounding

### v2

@@ 3.1.3 Slot-Based Token Compression
The slot-based token compression is inspired by the slot mechanism in SoftMoE (Puigcerver et al., 2023). However, in SoftMoE, the slots are not used to reduce sequence length, which is distinct from our primary objective. Moreover, some existing multi-modal LLMs compress the token number through cross-attention (Zhang et al., 2023; Bai et al., 2023; Ren et al., 2023). Our approach involves training only one matrix $\mathbf{\Phi}$, making it more computationally efficient and less data-consuming, and achieve better performance (Table 2).
3.2 VTG-IT-120K: Formatted Time-Sensitive Instruction Tuning Dataset
In this section, we introduce the VTG-IT-120K, which consists of 51.9K self-collected annotations and 67.6K annotations obtained from TimeIT (Ren et al., 2023). As shown in Figure 2(a), VTG-IT-120K primarily covers four distinct video temporal grounding tasks, including:
  • Moment Retrieval (63.2K): For the moment retrieval task, we use HiRESTgrounding (Zala et al., 2023a), QuerYD (Oncescu et al., 2021), DiDeMo (Hendricks et al., 2018b), and self-collected VTG-IT-MR.
  • Dense Video Captioning (37.2K): For the dense video captioning task, we use HiRESTstep (Zala et al., 2023a), COIN (Tang et al., 2019), ActivityNet Captions (Fabian Caba Heilbron & Niebles, 2015), and self-collected VTG-IT-DVC.
  • Video Summarization (15.2K): For the video summarization task, we use TVSum (Song et al., 2015), SumMe (Gygli et al., 2014), and self-collected VTG-IT-VS.
  • Video Highlight Detection (3.9K): For the video highlight detection task, we use self-collected VTG-IT-VHD.
The VTG-IT-X annotations for the four tasks are self-collected using 16K videos from the YT-Temporal-180M dataset (Zellers et al., 2021).
Data formatting.
All the tasks are formatted to QA pairs, and we use diverse questions to help models understand human intention. The answers are formatted for better transferring knowledge between different tasks. In detail, we reformat the annotations of different tasks, and use more similar answer formats among tasks. More detailed examples can be found in Appendix B.
  • Moment retrieval and dense video captioning. The GPT responses for these two tasks are formatted as ’start time - end time, description’. The moment retrieval task will consist of a single sentence, while the dense video captioning task may include multiple sentences.
  • Video summarization and video highlight detection. The responses for these two tasks are formatted as ’timestamp, score, description’. For the video highlight detection task, the description is the same as the query; for the video summarization task, the descriptions use placeholders like ’the highlight of the video’.
4 Experiments
4.1 Experiment Settings
Models and training configuration.
Unless otherwise specified, we employ ViT-G/14 from EVA-CLIP (Sun et al., 2023) and the Qformer from InstructBLIP (Dai et al., 2024) as the visual backbone. The vision projection layer and the weights of sequence embedding are initialized using the Video-LLaMA checkpoint (Zhang et al., 2023). For the language model backbone, we utilize LLaMA-2-7B (Touvron et al., 2023). The number of slots is set to 256. The models are trained on the VTG-IT-120k and a randomly sampled subset (97k) from the Valley dataset (Luo et al., 2023b). We randomly sample 96 frames for each video. Training is carried out using 16 ATN 910B NPUs with a batch size of 64, taking approximately 40 hours to complete. The fine-tuning also uses 16 NPUs and set the batch size to 16. We set the learning rate to 3e-5 initially, and train the models for 10 epochs by default.
Evaluation datasets, metrics, and baseline models.
We evaluate the model performance on three different tasks:
  • dense video captioning. We employ Youcook2 (Zhou et al., 2018) as the test dataset, and following the evaluation settings in TimeChat (Ren et al., 2023). Metrics including CIDEr (Vedantam et al., 2015), which assessing the quality of the captions; (2) SODA_c (Fujita et al., 2020) for story-level evaluation; (3) F1 score to measure the model’s ability to accurately locate timestamps;
  • Moment retrieval. We utilize test set of Charades-STA (Gao et al., 2017) for the moment retrieval task and report the recall at IOU thresholds of 0.5 and 0.7.
  • Video highlight detection. We employ the validation set of the QVHighlights dataset (Lei et al., 2021) and report the mean average precision (mAP) with IOU thresholds of 0.5 and 0.75, as well as the HIT@1, which represents the hit ratio of the highest scored clip.
For baseline models, we select Valley (Luo et al., 2023b), VideoChat (Li et al., 2023b), Video-ChatGPT (Maaz et al., 2023), and Video-LLaMA (Zhang et al., 2023) as examples of traditional video LLMs. For video LLMs specifically designed for VTG tasks, we choose TimeChat (Ren et al., 2023), VTimeLLM (Huang et al., 2023), Momentor (Qian et al., 2024), and HawkEye (Wang et al., 2024b)  [^1] Although related, LITA (Huang et al., 2024) employs self-built evaluation sets, making fair comparison difficult..
4.2 Results
Zero-shot performance of the VTG-LLM on various VTG tasks.
In Table 1, we evaluate the performance of the VTG-LLM on three VTG tasks, including dense video captioning (Youcook2), moment retrieval (Charades-STA), and video highlight detection (QVHighlights). The results indicate that: (1) The VTG-LLM model outperforms all existing 7B video LLM models on the zero-shot setting across all three datasets; (2) The performance of the VTG-LLM on the Charades-STA dataset for IOU=0.7 surpasses that of the VTimeLLM-13B model, indicating the superior performance of the VTG-LLM in accurately locating timestamps. (3) The improvement of the VTG-LLM on metrics that evaluate timestamp location accuracy, such as the F1 score for the Youcook2 dataset and Recall for the Charades-STA dataset, is more pronounced. This suggests that the performance gain of the VTG-LLM primarily stems from more accurate timestamp location. Additionally, performance on other metrics like CIDEr and SODA_c also improves.
Table 1: Zero-shot performance of algorithms over various tasks. We evaluated the performance of VTG-LLM using the Youcook2, Charades-STA, and QVHighlights datasets. We highlight the best results for each block using bold font. The Valley, VideoChat-Embed, Video-ChatGPT, and Video-LLaMA results are elaborated from previous studies (Ren et al., 2023; Huang et al., 2023; Qian et al., 2024). The results with gray color indicates unfair comparison (13B). The VTG-LLM samples 96 frames during training and testing stages.
@@ 4.2 Results
Quality Analysis.
In Figure 3, we compare the outputs of VTG-LLM and TimeChat in a zero-shot manner on the Youcook2 dataset. We observe that VTG-LLM’s outputs are more closely aligned with human intention. For instance, there are no repeated sentences with the same meaning.
Figure 3: Zero-shot case study. We compare the model outputs of VTG-LLM and TimeChat (Ren et al., 2023) on Youcook2 dataset in the zero-shot manner. The colored texts indicating repeated captions. More case studies can be found in Appendix C.
4.3 Ablation Studies
In this section, we aim to show that all the designed components are necessary for VTG-LLM to achieve the best results.
Table 2: Ablation studies on VTG-LLM. We perform ablation studies on the three components of the algorithm. For the Sequence-Time Embedding section, "SE Only" refers to using only sequence embedding. "TE Random Initialize" represents the random initialization of time embedding rather than the zero initialization we recommend. In the Time Token component, "No Time Token" signifies not using the time token, while "Time Token not Formatted" implies not formatting timestamps to the same length. "Token Embedding not Initialized" indicates randomly initialize the token embedding and LLM prediction head for time tokens. Lastly, for the Token Compression aspect, we compare slot-based token compression to several straightforward baselines. We highlight the best results of ablation studies using underline.
@@ 4.3 Ablation Studies
Ablation studies on types of token compression methods.
In the "Ablation Studies on Token Compression" section of Table 2, we present the performance of VTG-LLM using different token compression methods and include some naive baselines. The "Entropy Sampling" method involves sampling the number of tokens to 256 by selecting tokens with the maximum entropy, where the entropy is estimated using k-nearest neighbor distances (Kozachenko & Leonenko, 1987). The "Diverse Sampling" method involves sampling tokens to 256 using k-means++ (Arthur et al., 2007) to select tokens with the maximum distance. The "Cross Attention" method involves using a cross-attention layer to compress the number of tokens. The results indicate that: (1) The slot-based sampling achieves better performance among all compression methods; (2) Sampling-based methods perform poor compared to slot-based token compression method. We conjecture that this is because the sampling methods drop too much necessary information; (3) The cross-attention method performs even worse than sampling-based methods, which might be attributed to the attention mechanism requiring a large amount of data to perform well.
Ablation studies on sequence-time embedding.
In the "Ablation Studies on STE" section of Table 2, we show the performance of VTG-LLM when only using sequence embedding ("SE Only") and using sequence-time embedding but randomly initializing the weights of time-embedding ("TE Random Initialize"). The results indicate that: (1) the time-embedding is necessary for VTG-LLM to achieve the best result. Without the time-embedding, the performance of VTG-LLM on all datasets decreases; (2) Randomly initializing the time embedding significantly hurts the performance of VTG-LLM, making its performance worse than "SE Only".
Ablation studies on time-token.
In the "Ablation Studies on Time Token" section of Table 2, we show the performance of VTG-LLM when not using special time tokens ("No Time Token"), when not formatting the length of timestamps to six time tokens ("Time Token not Formatted"), and when do not initialize the token embedding and LLM head for time tokens ("Token Embedding not Initialized"). The results indicate that
@@
  • We found that the "No Time Token" setting tends to predict fewer timestamps on QVHighlights dataset, which may lead to better HIT@1 performance but a worse mAP score.
  • Formatting the timestamps using the same number of time tokens significantly enhances the model performance ("Time Token not Formatted"). We believe that this is because the formatted setting simplifies the task, making the models needless to identify the number of tokens in timestamps.
@@
Table 3: Fine-tuned performance of algorithms over various tasks. We fine-tune the Algorithm on Youcook2, Charades-STA, and QVHighlights datasets with a batch size of 16. We use gray color for traditional non-generalist models and task-specific models. Vid2Seq (Yang et al., 2023) fully training the LLM models and using 1B video data for pretraining.
Model Youcook2 Charades-STA QVHighlights
[…]
Non-Generative Models
InternVideo2-6B (Wang et al., 2024a) 70.0 49.0
VDI (Luo et al., 2023a) 52.3 31.4
Moment-DETR (Lei et al., 2021) 55.7 34.2 37.4 60.2
HL-CLIP (Han et al., 2024) 41.9 70.6
Task-Specific Models
Vid2Seq (Yang et al., 2023) 7.9 47.1 27.3
Vid2Seq (Visual Only) (Yang et al., 2023) 5.7 25.3 23.5
HawkEye 58.3 28.8
Generalist Models
TimeChat 3.4 11.0 19.5 46.7 23.7 21.7 37.9
VTG-LLM 3.6 13.4 20.6 57.2 33.4 24.1 41.3
Fine-tuned model performance of VTG-LLM.
In Table 3, we fine-tune the VTG-LLM models on Youcook2, Charades-STA, and QVHighlights datasets. From the results, we make the following observations: (1) The performance of VTG-LLM is significantly better than other generalist models across all three tasks; (2) The performance of VTG-LLM on the moment retrieval task (Charades-STA) is comparable to task-specific models like HawkEye, as well as non-generative models like VDI and Moment-DETR. (3) The task-specific model Vid2Seq (Yang et al., 2023) achieves superior performance on the Youcook2 dataset by fully training the LLM models and using 1B video data for pretraining. In contrast, we only fine-tune the LLM using LoRA and employ 120K instruction tuning data.
5 Conclusion, Limitation, and Future Works
@@
In this paper, we introduce VTG-IT-120K, a high-quality instruction tuning dataset that covers four video temporal grounding (VTG) tasks, including moment retrieval, dense video captioning, video highlight detection, and video summarization. Furthermore, we propose the VTG-LLM, an improved video LLM model for integrating knowledge about timestamps, enhancing the model performance on VTG tasks. Extensive numerical results demonstrate the superior zero-shot performance of VTG-LLM over state-of-the-art video LLM models on Charades-STA, QVHighlights, and Youcook2 datasets, with each proposed component contributing to individual performance gains.
However, there are still some constraints with the VTG-LLM that warrant further research. For instance, the VTG-LLM’s capacity for dense video captioning and highlight detection remains subpar compared to non-generative or task-specific models. Therefore, exploring more advanced methods to rectify this issue is a worthwhile endeavor. Additionally, while this study primarily focuses on the accuracy of timestamp location, it would also be beneficial to investigate techniques to enhance other aspects, such as the precision of salient scores and the quality of captions. Moreover, we did not incorporate the audio elements of the videos in this study. Integrating these in the future could potentially enhance the overall performance.
Appendix A Experiments
A.1 Detailed Experiment Settings
In this section, we present the detailed experimental settings used in our study. We employ ViT-G/14 from EVA-CLIP (Sun et al., 2023) and Qformer from InstructBLIP (Dai et al., 2024) as the visual backbone. The number of queries for Qformer is set to 32. The vision projection layer and the weights of the sequence embedding are initialized using the Video-LLaMA checkpoint (Zhang et al., 2023). To address the mismatched size of the sequence embedding, we use interpolation to extend the size from 32 to 96. The time-embedding (8192) is initialized with zeros. We set the number of slots to 256.
For the language model backbone, we utilize LLaMA-2-7B (Touvron et al., 2023). The LoRA training is conducted on "q_proj", "k_proj", "v_proj", and "o_proj". Special time tokens are added to the LLaMA tokenizer, and we set the prediction head of LLM and the weights of token embedding to be trainable. The maximum text length is set to 2048.
The models are trained on the VTG-IT-120k and a randomly sampled subset (97k) from the Valley dataset (Luo et al., 2023b). We randomly sample 96 frames for each video. Training is carried out using 16 ATN 910B GPUs with a batch size of 64 and takes approximately 40 hours to complete. The training employs DDP and requires about 30GB of GPU storage for each GPU. Fine-tuning also uses 16 GPUs and sets the batch size to 16. We fine-tune the VTG-LLM model for 10 epochs on Charades dataset, 16 epochs for Youcook2 dataset, and 20 epochs for QVHighlights dataset. We set the initial learning rate to 3e-5 and train the models for 10 epochs by default. The weight decay is set to 0.05, and the warm-up steps use $0.6E$, where $E$ represents the number of iterations for each epoch.
Appendix B Data Annotation Examples
We present annotation examples for moment retrieval (Figure 4), dense video captioning (Figure 5), video highlight detection (Figure 6), and video summarization (Figure 7). It should be noted that in the VTG-IT-VHD dataset, only clips with high scores are annotated. This leads to a relatively consistent score distribution among the annotated clips. As a result, the VTG-IT-VHD dataset assists models in locating relevant video clips; however, the significant scores may not be accurately predicted in a zero-shot manner due to differing scoring criteria. This issue can be addressed by fine-tuning the models on target datasets, which helps the models understand the specific scoring criteria.
Figure 4: Annotation example of moment retrieval task.
Figure 5: Annotation example of dense video captioning task.
Figure 6: Annotation example of dense video highlight detection task. It should be noted that in the VTG-IT-VHD dataset, only clips with high scores are annotated. This leads to a relatively consistent score distribution among the annotated clips.
Figure 7: Annotation example of dense video summarization task.
Appendix C Additional Case Studies
We show the additional zero-shot case studies on dense video captioning (Figure 8), moment retrieval (Figure 9), and video highlight detection (Figure 10) tasks.
Figure 8: Zero-shot case study of dense video captioning task.
Figure 9: Zero-shot case study of moment retrieval task.
Figure 10: Zero-shot case study of video highlight detection task.

### v3

@@ Introduction
The slot-based token compression is inspired by the slot mechanism in SoftMoE (Puigcerver et al. 2023). However, in SoftMoE, the slots are not used to reduce sequence length, which is distinct from our primary objective. Moreover, some existing multi-modal LLMs compress the token number through cross-attention (Zhang, Li, and Bing 2023; Bai et al. 2023; Ren et al. 2023). Our approach involves training only one matrix $\mathbf{\Phi}$, making it more computationally efficient and less data-consuming, and thereby achieve better performance (Table 2) on relatively small-scale instruction tuning datasets.
@@
Table 1: Zero-shot performance of algorithms over various tasks.
VTG-IT-120K: Formatted Time-Sensitive Instruction Tuning Dataset
In this section, we introduce VTG-IT-120K, a dataset comprising 120K publicly available video-text pairs, building on the TimeIT dataset (Ren et al. 2023). Additionally, we have re-annotated 51.9K low-quality video annotations using Gemini 1.5-Pro  [^1] The details of the annotation process can be found in Appendix (Guo et al. 2024).. A comparison between the original and new annotations can be seen in Figure 3. It is clear that the revised captions are more succinct and contain less information unrelated to the visual content. The VTG-IT-120K dataset encompasses four distinct video temporal grounding tasks.
  • Moment Retrieval (63.2K): For the moment retrieval task, we use HiRESTgrounding (Zala et al. 2023a), QuerYD (Oncescu et al. 2021), DiDeMo (Hendricks et al. 2018), and VTG-IT-MR.
  • Dense Video Captioning (37.2K): For the dense video captioning task, we use HiRESTstep (Zala et al. 2023a), COIN (Tang et al. 2019), ActivityNet Captions (Fabian Caba Heilbron and Niebles 2015), and VTG-IT-DVC.
  • Video Summarization (15.2K): For the video summarization task, we use TVSum (Song et al. 2015), SumMe (Gygli et al. 2014), and VTG-IT-VS.
  • Video Highlight Detection (3.9K): For the video highlight detection task, we use VTG-IT-VHD.
The VTG-IT-X annotations for the four tasks are re-annotated using 16K videos from the YT-Temporal-180M dataset (Zellers et al. 2021).
Data formatting.
All tasks are structured as QA pairs, with varied questions designed to help models comprehend human intent. The answers are formatted to facilitate knowledge transfer between different tasks. Detailed examples are provided in the Appendix (Guo et al. 2024).
Experiments
In the experimental section, we primarily aim to address the following questions  [^2] Fine-tuned performance, results on ActivityNet Captions, more ablation studies, qualitative comparisons, and case studies can be found in Appendix (Guo et al. 2024).:
  • Q1. Will VTG-LLM achieve better zero-shot performance compared to SOTA video LLMs?
  • Q2. Is it necessary to use time embeddings, and what will happen if time embeddings are randomly initialized?
  • Q3. How will the special time tokens affect the model’s performance?
  • Q4. Will the slot-based compression method outperform other compression methods?
Figure 3: Example of new annotations.
Experiment Settings
Models and training configuration.
Unless otherwise specified, we employ ViT-G/14 from EVA-CLIP (Sun et al. 2023) and the Qformer from InstructBLIP (Dai et al. 2024) as the visual backbone. The vision projection layer and the weights of sequence embedding are initialized using the Video-LLaMA checkpoint (Zhang, Li, and Bing 2023). For the language model backbone, we utilize LLaMA-2-7B (Touvron et al. 2023). The number of slots is set to 256. The models are trained on the VTG-IT-120k and a randomly sampled subset (97k) from the Valley dataset (Luo et al. 2023b). For the training process, we initially divide the videos uniformly into 96 segments, from each of which we then randomly select one frame. For the testing process, we uniformly select 96 frames from the entire video. Training is carried out using 16 ATN 910B NPUs with a batch size of 64. The fine-tuning also uses 16 NPUs and set the batch size to 16. We set the learning rate to 3e-5 initially, and train the models for 10 epochs by default.
Evaluation datasets, metrics, and baseline models.
We evaluate the model performance on three different tasks:
  • dense video captioning. We employ Youcook2 (Zhou, Xu, and Corso 2018) as the test dataset, and following the evaluation settings in TimeChat (Ren et al. 2023). Metrics including CIDEr (Vedantam, Lawrence Zitnick, and Parikh 2015), which assessing the quality of the captions; (2) SODA_c (Fujita et al. 2020) for story-level evaluation; (3) F1 score to measure the model’s ability to accurately locate timestamps;
  • Moment retrieval. We utilize test set of Charades-STA (Gao et al. 2017) for the moment retrieval task and report the recall at IOU thresholds of 0.5 and 0.7.
  • Video highlight detection. We employ the validation set of the QVHighlights dataset (Lei, Berg, and Bansal 2021) and report the mean average precision (mAP) with IOU thresholds of 0.5 and 0.75, as well as the HIT@1, which represents the hit ratio of the highest scored clip.
For baseline models, we select Valley (Luo et al. 2023b), VideoChat (Li et al. 2023a), Video-ChatGPT (Maaz et al. 2023), and Video-LLaMA (Zhang, Li, and Bing 2023) as examples of traditional video LLMs. For video LLMs specifically designed for VTG tasks, we choose TimeChat (Ren et al. 2023), VTimeLLM (Huang et al. 2023), Momentor (Qian et al. 2024), and HawkEye (Wang et al. 2024b)  [^3] Although related, LITA (Huang et al. 2024) employs self-built evaluation sets, making fair comparison difficult..
Numerical Results
R1. VTG-LLM achieves superior zero-shot performance on various VTG tasks.
In Table 1, we evaluate the performance of the VTG-LLM on three VTG tasks, including dense video captioning (Youcook2), moment retrieval (Charades-STA), and video highlight detection (QVHighlights). The results indicate that: (1) The VTG-LLM model outperforms all existing 7B video LLM models on the zero-shot setting across all three datasets; (2) The improvement of the VTG-LLM on metrics that evaluate timestamp location accuracy, such as the F1 score for the Youcook2 dataset and Recall for the Charades-STA dataset, is more pronounced. This suggests that the performance gain of the VTG-LLM primarily stems from more accurate timestamp location. Additionally, performance on other metrics like CIDEr and SODA_c also improves. (3) As shown in Table 3, the performance of VTG-LLM remains relatively robust across different video durations.
@@
Table 2: Ablation studies on VTG-LLM.
YouCook2 CIDEr METEOR F1 SODA_c
$[0s,180s)$ 6.3 2.3 20.0 1.9
$[180s,240s)$ 4.6 1.8 18.8 1.4
$[240s,300s)$ 5.0 1.9 16.6 1.6
$[300s,420s)$ 5.6 2.1 17.9 1.5
$[420s,inf)$ 4.1 1.6 15.1 1.3
Table 3: Performance with different video durations.
R2.Time embedding boost the model performance when using zero initialization technique.
In the ”Ablation Studies on STE” section of Table 2, we show the performance of VTG-LLM when only using sequence embedding (”SE Only”) and using sequence-time embedding but randomly initializing the weights of time-embedding (”TE Random Initialize”). The results indicate that: (1) the time-embedding is necessary for VTG-LLM to achieve the best result. Without the time-embedding, the performance of VTG-LLM on all datasets decreases; (2) Randomly initializing the time embedding significantly hurts the performance of VTG-LLM.
R3. Special time tokens significantly boost temporal grounding capacity with the cost of slightly caption quality reduction.
In the ”Ablation Studies on Time Token” section of Table 2, we show the performance of VTG-LLM when not using special time tokens (”No Time Token”), when not formatting the length of timestamps to six time tokens (”Time Token not Formatted”), and when do not initialize the token embedding and LLM head for time tokens (”Token Embedding not Initialized”). The results indicate that
@@
  • We found that the ”No Time Token” setting tends to predict fewer timestamps on QVHighlights dataset, which may lead to better HIT@1 performance but a worse mAP score.
  • Formatting the timestamps using the same number of time tokens significantly enhances the model performance (”Time Token not Formatted”). We believe that this is because the formatted setting simplifies the task, making the models needless to identify the number of tokens in timestamps.
@@
R4. Slot-based token compression outperform naive compression baselines.
In the ”Ablation Studies on Token Compression” section of Table 2, we present the performance of VTG-LLM using different token compression methods and include some naive baselines. The ”Entropy Sampling” method involves sampling the number of tokens to 256 by selecting tokens with the maximum entropy, where the entropy is estimated using k-nearest neighbor distances (Kozachenko and Leonenko 1987). The ”Diverse Sampling” method involves sampling tokens to 256 using k-means++ (Arthur, Vassilvitskii et al. 2007) to select tokens with the maximum distance. The ”Cross Attention” method involves using a cross-attention layer to compress the number of tokens. The results indicate that: (1) The slot-based sampling achieves better performance among all compression methods; (2) Sampling-based methods perform poor compared to slot-based token compression method. We conjecture that this is because the sampling methods drop too much necessary information; (3) The cross-attention method performs even worse than sampling-based methods, which might be attributed to the attention mechanism requiring a large amount of data to perform well.
Conclusion, Limitation, and Future Works
@@
In this paper, we propose the VTG-LLM, an improved video LLM model for integrating knowledge about timestamps, enhancing the zero shot performance on four downstream VTG tasks. Extensive numerical results demonstrate the superior zero-shot performance of VTG-LLM over state-of-the-art video LLM models on Charades-STA, QVHighlights, and Youcook2 datasets, with each proposed component contributing to individual performance gains.
Nonetheless, there are certain limitations in the VTG-LLM that necessitate additional research. For example, while this study mainly concentrates on the accuracy of timestamp locations, it would be advantageous to explore methods for improving other aspects, such as the precision of salient scores and the quality of captions. Furthermore, we did not incorporate the audio components of the videos in this study. Moreover, evaluating the performance of slot-based compression on a more diverse set of tasks is also worth investigating.
Acknowledgments
This work is supported in part by the funding from Shenzhen Institute of Artificial Intelligence and Robotics for Society, in part by the Shenzhen Key Lab of Crowd Intelligence Empowered Low-Carbon Energy Network (Grant No. ZDSYS20220606100601002) , in part by Shenzhen Stability Science Program 2023, and in part by the Guangdong Provincial Key Laboratory of Future Networks of Intelligence (Grant No. 2022B1212010001).